# Partially Stochastic Infinitely Deep Bayesian Neural Networks

### v1

@@ A.1 Proof of Lemma 1
(Theorem 1 by Sharma et al. 2023). Let $X$ be a random variable taking values in $\mathcal{X}$, where $\mathcal{X}$ is a compact subspace of $\mathbb{R}^{d}$, and let $Y$ be a random variable taking values in $\mathcal{Y}$, where $\mathcal{Y}\subseteq\mathbb{R}^{n}$. Further, let $f_{\theta}:\mathbb{R}^{m}\times\mathcal{X}\rightarrow\mathcal{Y}$ represent a neural network architecture with a universal approximation property with deterministic parameters $\theta\in\Theta$, such that, for input $X=x$, the network produces outputs $f_{\theta}(Z,x)$, where $Z=\left\{Z_{1},\ldots,Z_{m}\right\},Z_{i}\in\mathbb{R}$, are the random variables in the network, which are Gaussian, independent of $X$, and have finite mean and variance. If there exists a continuous generator function, $\tilde{f}:\mathbb{R}^{m}\times$ $\mathcal{X}\rightarrow\mathcal{Y}$, for the conditional distribution $Y\mid X$, then $f_{\theta}$ can approximate $Y\mid X$ arbitrarily well. Formally, $\forall\varepsilon>0,\lambda<$infty,
@@
Appendix B Uncertainty Quantification - ROC Curves
@@ Appendix C Further Training Visualisations
Appendix D Experimental setting

### v2

@@ Appendix A Proofs
(Theorem 1 by Sharma et al. (2023)). Let $X$ be a random variable taking values in $\mathcal{X}$, where $\mathcal{X}$ is a compact subspace of $\mathbb{R}^{d}$, and let $Y$ be a random variable taking values in $\mathcal{Y}$, where $\mathcal{Y}\subseteq\mathbb{R}^{n}$. Further, let $f_{\theta}:\mathbb{R}^{m}\times\mathcal{X}\rightarrow\mathcal{Y}$ represent a neural network architecture with a universal approximation property with deterministic parameters $\theta\in\Theta$, such that, for input $X=x$, the network produces outputs $f_{\theta}(Z,x)$, where $Z=\left\{Z_{1},\ldots,Z_{m}\right\},Z_{i}\in\mathbb{R}$, are the random variables in the network, which are Gaussian, independent of $X$, and have finite mean and variance. If there exists a continuous generator function, $\tilde{f}:\mathbb{R}^{m}\times$ $\mathcal{X}\rightarrow\mathcal{Y}$, for the conditional distribution $Y\mid X$, then $f_{\theta}$ can approximate $Y\mid X$ arbitrarily well. Formally, $\forall\varepsilon>0,\lambda<$infty,
@@
Appendix B More on Uncertainty Quantification
Refer to captionRefer to captionRefer to captionRefer to caption
@@ Appendix C Further Training Visualisations
Refer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to caption
@@
Refer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to caption
@@ Appendix D Further Ablations
D.1 Computational Complexity Overview
We will now provide a breakdown of the time complexities for each model. Let $D_{h},D_{w}$ denote the dimension of the hidden state $h_{t}$ and weights $w_{t}$ respectively, and let $D=D_{h}+D_{w}$. At inference, we have:
  • Shared Complexity in deterministic part (i.e. induced by $\Delta t$ in equation (10)): Both our and (Xu et al., 2022)’s approaches share the complexity $\mathcal{O}(F(D)T)$ for the deterministic part of the Euler-Maruyama discretization, where $T$ is the total number of time steps and $F(D)$ represents the cost of evaluating the functions $f_{g}$ and $f_{h}$ at each time step.
  • SDE Part Complexity (i.e. induced by $dB_{t}$ in equation (10)) - Original SDE-BNN Model: The SDE-BNN introduces additional complexity proportional to $D_{w}T$, corresponding to an additional $D_{w}$-dimensional Gaussian random variable to simulate at each time step (stochastic part of the Euler-Maruyama discretization)
  • SDE Part Complexity - PSDE-BNN Model
    • Vertical Cut: Introducing vertical cut partitions the time into deterministic and stochastic phases, reducing the SDE part complexity to $\mathcal{O}((D_{w}T_{\text{subset}})$ where $T_{\text{subset}}$ is the stochastic phase duration. We can write this as $\mathcal{O}(rD_{w}T)$, given a stochastic ratio $r=\frac{T_{\text{subset}}}{T}\leq 1$.
    • Horizontal Cut: This model limits the stochastic dimensions to $D_{s}<D_{w}$, reducing the SDE part complexity to $\mathcal{O}(D_{s}T)$ (or $\mathcal{O}(rD_{s}T)$ if combined with a vertical cut).
We can then summarize the total time complexities in the following way:
Model Time Complexity
SDE-BNN $\mathcal{O}((F(D)+D_{w})T)$
PSDE-BNN with Vertical cut $\mathcal{O}((F(D)+D_{w})T)$
PSDE-BNN with Horizontal cut $\mathcal{O}((F(D)+D_{s})T)$
PSDE-BNN with Vertical and Horizontal cuts $\mathcal{O}((F(D)+rD_{s})T)$
where $r<1$ and $D_{s}<D_{w}$.
The distinction between $D_{w}$ and $D_{h}$ is key, as $D_{h}$ is typically much larger, making stochastic simulations notably more demanding. Thus, by reducing the stochastic portion, our models significantly enhance efficiency without compromising the depth of stochastic analysis.
Appendix E Experimental setting